Can someone help me solve this problem?
I need to solve for x and check my result:
-4(2x-3) = -8x + 5
Thanks,
Lyn

Answer by checkley71(8403)

-4(2X+3)=-8X+5 OR -8X-12=-8X+5 OR -8X+8X=5+12 OR OX=17 OR NO SOLUTIONS FOR X